# Thermodynamics of ideal spin fluids and pseudo-gauge ambiguity

### v1

@@ Appendix A Useful identities in thermodynamic equilibrium
In this appendix, we collect some useful identities for spin fluids in thermodynamic equilibrium for later calculations. Firstly, note that the thermodynamic derivatives in the basis of scalars $\{T,\alpha^{2},w^{2},\alpha\cdot w,\nu\}$ are related to the ones in the basis $\{T,a^{2},w^{2},a\cdot\omega,\mu\}$ as
@@
In these appendices, we will primarily use non-scaled variables to avoid confusion, but will make use of $\mathrm{d}_{T}$ and $\mathrm{d}_{A}$ thermodynamic derivatives.
@@
where we have defined $\bar{\epsilon}^{\mu\nu\rho}=u_{\lambda}\epsilon^{\lambda\mu\nu\rho}$. We also note that
@@
$\partial_{\mu}f$ $=-a_{\mu}T\mathrm{d}_{T}f-\mathopen{}\mathclose{{\left(\omega^{2}a_{\mu}-(a\cdot\omega)w_{\mu}}}\right)\frac{\mathrm{d}_{A}f}{T^{2}}.$ (39)
@@
These identities are pseudo-gauge-independent, i.e. they apply in any pseudo-gauge. Also, note that the first two identities completely fix the $\mathcal{E}$ and $\mathcal{Q}_{\ell}$ components in terms of the others. The third identity completely fixes $\mathcal{Q}_{a}$, while the fourth identity fixes $\mathcal{Q}_{\omega}$ up to a term $\sim T^{3}f(\nu,\alpha^{2},\omega^{2},a\cdot\omega)$. Similarly, there is one identity arising from the conservation of charge current
@@
These can be used to fix the $\mathcal{N}_{\omega}$ and $\mathcal{S}_{\omega}$ components up to terms $\sim T^{2}f(\nu,\alpha^{2},\omega^{2},a\cdot\omega)$.
Appendix B Independent tensor structures for pseudo-gauge ambiguity
@@
Appendix C Details of improvement terms for charge current
@@
We can verify that $\delta\mathcal{N}_{a}$ and $\delta\mathcal{N}_{\omega}$ satisfy the charge conservation identity in eq.˜42.
@@
Appendix D Details of improvement terms for Belinfante energy-momentum tensor
@@
One may check that these expressions satisfy the conservation identities in appendix˜A.
@@
Appendix E Details of thermodynamic constraints and thermodynamic relations
@@
After substituting the pseudo-gauge improvements from appendices˜C and 54, these yield PDEs for 10 pseudo-gauge parameters: $\gamma_{4}$, $\gamma_{17}$, $\gamma_{19}$, $\gamma_{21}$, $\gamma_{23}$, $\gamma_{33}$, $\gamma_{34}$, $\gamma_{36}$, $\lambda_{6}$, and $\lambda^{S}_{6}$. In all the microscopic examples we discuss, the respective components of the currents are already zero. Therefore, we can trivially solve the constraints in eq.˜56 by setting the relevant pseudo-gauge parameters to zero. Of course, these parameters may be more non-trivial to solve for in generic microscopic theories. The remaining constraints are given as
@@
which provide PDEs to solve for 20 pseudo-gauge parameters: $\gamma_{1}$, $\gamma_{2}$, $\gamma_{7}$, $\gamma_{8}$, $\gamma_{9}$, $\gamma_{13}$, $\gamma_{15}$, $\gamma_{16}$, $\gamma_{25}$, $\gamma_{27}$, $\gamma_{29}$, $\gamma_{31}$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{4}$, $\lambda_{5}$, $\lambda_{1}^{S}$, $\lambda_{2}^{S}$, $\lambda_{4}^{S}$, and $\lambda_{5}^{S}$. Using the thermodynamic relations on the right-hand side, we find that two combinations of these reduce to the first two conservation identities in appendix˜A. The third equation becomes the formula for EoS in eq.˜22, while the remaining combinations yield the thermodynamic constraints outlined in eq.˜21.
@@
Note that the constraint on $\mathcal{N}^{\prime}$ splits into two constraints for $\partial\mathcal{N}^{\prime}/\partial a^{2}$ and $\partial\mathcal{N}^{\prime}/\partial\omega^{2}$, and similarly for $\mathcal{S}^{\prime}$. At leading order the pseudo-gauge improvements in eqs.˜53 and C reduce to
@@
Substituting these back into eq.˜59, we find the PDEs
@@
Appendix F Details of non-interacting Dirac fermions
@@
Note that the authors in Palermo et al. [2021], Becattini et al. [2021], Buzzegoli et al. [2017], Becattini and Singh [2025] are using the conventions $\eta_{\mu\nu}=\operatorname{diag}(1,-1,-1,-1)$ and $\epsilon^{0123}=1$, as opposed to our conventions of $\eta_{\mu\nu}=\operatorname{diag}(-1,1,1,1)$ and $\epsilon_{0123}=1$. We have presented the pseudo-gauge parameters needed to cast appendix˜F into thermodynamic form in eq.˜28. The ensuing equation of state is given in eq.˜29. More details on the derivation of these results can be found in the supplementary Mathematica notebook.
@@
$\mathcal{N}$ $=\frac{2m^{3}}{\pi^{2}}\sum_{n=1}^{\infty}(-1)^{n+1}\sinh(\nu n)\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(1+\frac{n^{2}\omega^{2}}{8T^{2}}}}\right)\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}^{2}}+\frac{n^{2}a^{2}}{24T^{2}}\mathopen{}\mathclose{{\left(\frac{K_{3}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}}-\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}^{2}}}}\right)}}\right)+\mathcal{O}(\mu_{\mu\nu}^{4}),$
@@
Appendix G Details of non-interacting scalar fields
@@
where $\xi$ is an arbitrary parameter arising from coupling to curved spacetime Becattini et al. [2021]. A few comments are in order. Massless scalar fields only admit equilibrium states for $\mu=0$. Hence, we can only read off the $\mu\to 0$ part of appendix˜G by comparing with the currents in equilibrium. We have reinstated the $\mu\neq 0$ part of the constitutive relations by comparing with the results for massive scalar fields in the $m\to 0$ limit Buzzegoli et al. [2017]; see below. As such, appendix˜G should be interpreted in the hydrodynamic sense, valid for small perturbations near equilibrium.
@@
The parameters of pseudo-gauge improvements to cast appendix˜G into conformal thermodynamic form are given as
@@
$\mathcal{P}$ $=\frac{m^{4}}{\pi^{2}}\sum_{n=1}^{\infty}\cosh(\nu n)\Bigg(\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}^{2}}+\frac{n^{2}a^{2}}{24T^{2}}\mathopen{}\mathclose{{\left(4(1-4\xi)\frac{\text{K2}(\text{T0})}{x_{n}^{2}}+(24\xi-5)\frac{K_{3}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}}}}\right)+\frac{\xi n^{2}\omega^{2}}{T^{2}}\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}^{2}}\Bigg)+\mathcal{O}(\mu_{\mu\nu}^{4}),$

### v2

@@ Appendix B Useful identities in thermodynamic equilibrium
In this appendix, we collect some useful identities for spin fluids in thermodynamic equilibrium for later calculations. Firstly, note that the thermodynamic derivatives in the basis of scalars $\{T,\alpha^{2},w^{2},\alpha\cdot w,\nu\}$ are related to the ones in the basis $\{T,a^{2},\omega^{2},a\cdot\omega,\mu\}$ as
@@
In these appendices, we will primarily use the non-scaled variables to avoid confusion, but will make use of $\mathrm{d}_{T}$ and $\mathrm{d}_{A}$ thermodynamic derivatives.
@@
where we have defined the spatial Levi-Civita symbol $\bar{\epsilon}^{\mu\nu\rho}=u_{\lambda}\epsilon^{\lambda\mu\nu\rho}$. We also note that
@@
$\partial_{\mu}f$ $=-a_{\mu}T\mathrm{d}_{T}f-\mathopen{}\mathclose{{\left(\omega^{2}a_{\mu}-(a\cdot\omega)\omega_{\mu}}}\right)\frac{\mathrm{d}_{A}f}{T^{2}}.$ (53)
@@
These identities are pseudo-gauge-independent, i.e. they apply in any pseudo-gauge. Also, note that the first two identities completely fix the $\mathcal{E}$ and $\mathcal{Q}_{\ell}$ components in terms of the others. The third identity completely fixes $\mathcal{Q}_{a}$, while the fourth identity fixes $\mathcal{Q}_{\omega}$ up to a term $\sim T^{3}f(\nu,\alpha^{2},w^{2},\alpha\cdot w)$. Similarly, there is one identity arising from the conservation of charge current
@@
These can be used to fix the $\mathcal{N}_{\omega}$ and $\mathcal{S}_{\omega}$ components up to terms $\sim T^{2}f(\nu,\alpha^{2},w^{2},\alpha\cdot w)$.
Appendix C Independent tensor structures for pseudo-gauge ambiguity
@@
Appendix D Details of improvement terms for charge current
@@
We can verify that $\delta\mathcal{N}_{a}$ and $\delta\mathcal{N}_{\omega}$ satisfy the charge conservation identity in eq.˜56.
@@
Appendix E Details of improvement terms for Belinfante energy-momentum tensor
@@
One may check that these expressions satisfy the conservation identities in appendix˜B.
@@
Appendix F Details of thermodynamic constraints and thermodynamic relations
@@
After substituting the pseudo-gauge improvements from appendices˜D and 68, these yield PDEs for 10 pseudo-gauge parameters: $\gamma_{4}$, $\gamma_{17}$, $\gamma_{19}$, $\gamma_{21}$, $\gamma_{23}$, $\gamma_{33}$, $\gamma_{34}$, $\gamma_{36}$, $\lambda_{6}$, and $\lambda^{S}_{6}$. In all the microscopic examples we discuss, the respective components of the currents are already zero. Therefore, we can trivially solve the constraints in eq.˜70 by setting the relevant pseudo-gauge parameters to zero. Of course, these parameters may be more non-trivial to solve for in generic microscopic theories. The remaining constraints are given as
@@
which provide PDEs to solve for 20 pseudo-gauge parameters: $\gamma_{1}$, $\gamma_{2}$, $\gamma_{7}$, $\gamma_{8}$, $\gamma_{9}$, $\gamma_{13}$, $\gamma_{15}$, $\gamma_{16}$, $\gamma_{25}$, $\gamma_{27}$, $\gamma_{29}$, $\gamma_{31}$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{4}$, $\lambda_{5}$, $\lambda_{1}^{S}$, $\lambda_{2}^{S}$, $\lambda_{4}^{S}$, and $\lambda_{5}^{S}$. Using the thermodynamic relations on the right-hand side, we find that two combinations of these reduce to the first two conservation identities in appendix˜B. The third equation becomes the formula for EoS in eq.˜22, while the remaining combinations yield the thermodynamic constraints outlined in eq.˜21.
@@
Note that the constraint on $\mathcal{N}^{\prime}$ splits into two constraints for $\partial\mathcal{N}^{\prime}/\partial a^{2}$ and $\partial\mathcal{N}^{\prime}/\partial\omega^{2}$, and similarly for $\mathcal{S}^{\prime}$. At leading order the pseudo-gauge improvements in eqs.˜67 and D reduce to
@@
Substituting these back into eq.˜73, we find the PDEs
@@
Appendix G Details of non-interacting Dirac fermions
@@
Note that the authors in Palermo et al. [2021], Becattini et al. [2021], Buzzegoli et al. [2017], Becattini and Singh [2025] are using the conventions $\eta_{\mu\nu}=\operatorname{diag}(1,-1,-1,-1)$ and $\epsilon^{0123}=1$, as opposed to our conventions of $\eta_{\mu\nu}=\operatorname{diag}(-1,1,1,1)$ and $\epsilon_{0123}=1$. We have presented the pseudo-gauge parameters needed to cast appendix˜G into the thermodynamic form in eq.˜28. The ensuing equation of state is given in eq.˜29. More details on the derivation of these results can be found in the supplementary Mathematica notebook.
@@
$\mathcal{N}$ $=\frac{2m^{3}}{\pi^{2}}\sum_{n=1}^{\infty}(-1)^{n+1}\sinh(\nu n)\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(1+\frac{n^{2}\omega^{2}}{8T^{2}}}}\right)\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}}+\frac{n^{2}a^{2}}{24T^{2}}\mathopen{}\mathclose{{\left(K_{3}\mathopen{}\mathclose{{\left(x_{n}}}\right)-\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}}}}\right)}}\right)+\mathcal{O}(\mu_{\mu\nu}^{4}),$
@@
Appendix H Details of non-interacting scalar fields
@@
where $\xi$ is an arbitrary parameter arising from coupling to curved spacetime Becattini et al. [2021]. A few comments are in order. Massless scalar fields only admit equilibrium states for $\mu=0$. Hence, we can only read off the $\mu\to 0$ part of appendix˜H by comparing with the currents in equilibrium. We have reinstated the $\mu\neq 0$ part of the constitutive relations by comparing with the results for massive scalar fields in the $m\to 0$ limit Buzzegoli et al. [2017]; see below. As such, appendix˜H should be interpreted in the hydrodynamic sense, valid for small perturbations near equilibrium.
@@
The parameters of pseudo-gauge improvements to cast appendix˜H into conformal thermodynamic form are given as
@@
$\mathcal{P}$ $=\frac{m^{4}}{\pi^{2}}\sum_{n=1}^{\infty}\cosh(\nu n)\Bigg(\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}^{2}}+\frac{n^{2}a^{2}}{24T^{2}}\mathopen{}\mathclose{{\left(12(1-4\xi)\frac{K_{2}(x_{n})}{x_{n}^{2}}+(24\xi-5)\frac{K_{3}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}}}}\right)+\frac{\xi n^{2}\omega^{2}}{T^{2}}\frac{K_{2}\mathopen{}\mathclose{{\left(x_{n}}}\right)}{x_{n}^{2}}\Bigg)+\mathcal{O}(\mu_{\mu\nu}^{4}),$